The smiths invested $7500, part a 8% simple interest and part at 5% simple interest. If they received a total annual interest of $525, how much did they invest at each rate?
Answer by palanisamy(496)

Let the amount invested by Smith in 8% simlple interest = x
Then, the amount invested by Smith in 5% simlple interest = 7500-x
Total interest is x*(8/100)+(7500-x)*(5/100)= 525
8x+37500-5x=52500
3x=52500-37500
x=15000/3
x=5000
So, the amount invested by Smith in 8% simlple interest = 5000
and the amount invested by Smith in 5% simlple interest = 7500-5000=2500